# Neutrino masses from new Weinberg-like operators: Phenomenology of TeV scalar multiplets

### v1

@@ Abstract
Majorana Neutrino Masses, Effective Field Theory, Electroweak Precision Tests, New Scalar multiplets, Pseudo-Nambu-Goldstone boson, Collider Searches, $W-$boson mass, Lepton number violation
@@ 1 Introduction
Considering that $C^{(i)}_{n}\equiv c^{(i)}_{n}/\Lambda$, with $c^{(i)}_{n}$ a dimensionless parameter and $\Lambda$ the new physics scale, Eq. (3) yields an explanation for the smallness of neutrino masses, which scale as $v^{2}/\Lambda$, where $v$ is the Vacuum Expectation Value (VEV) of the SM Higgs field, $\left\langle H\right\rangle\equiv v=174$ GeV. For order one couplings, $c^{(0)}_{5,a}\sim\mathcal{O}(1)$, reproducing the atmospheric mass scale, $m^{(0)}_{\nu}\gtrsim\mathcal{O}(0.05)$ eV, demands $\Lambda\lesssim 10^{14}$ GeV, close to the scale of Grand Unified Theories, this is the seesaw mechanism Yanagida:1980xy . If this is the origin of neutrino masses, it will therefore be very difficult to probe the new physics responsible for lepton number violation.
@@
The possible new Weinberg-like operators with up to two new scalars $\Phi_{1},\Phi_{2}$ are:
@@
In this work, we study and provide a catalogue of the new scenarios generated by these new Weinberg-like operators constructed with extra scalar multiplets that take a VEV. We consider natural scenarios, in which the new scalar multiplets obtain suppressed induced VEVs, i.e., $v_{i}\sim v^{3}/m_{\Phi}^{2}$. The most appealing scenarios are the ones in which the dominant contribution to neutrino masses is provided by the new operators. The UV models that generate those at tree level are considered genuine and will be studied in a companion paper us:2023two .
@@ 2.1 Neutrino masses in the Standard Model Effective Field Theory
$\mathcal{O}^{(0)}_{5,a}=(HL)_{\mathbf{1}}(HL)_{\mathbf{1}}\,,\quad$ $\mathcal{O}^{(0)}_{5,b}=(HL)_{\mathbf{3}}(HL)_{\mathbf{3}}$ (6)
$\mathcal{O}^{(0)}_{5,c}=(HH)_{\mathbf{3}}(LL)_{\mathbf{3}}\,,\quad$ $\mathcal{O}^{(0)}_{5,d}=(HH)_{\mathbf{1}}(LL)_{\mathbf{1}}.$
Since the singlet $(HH)_{1}=0$, the last operator identically vanishes. The other cases yield non-zero neutrino masses and their UV completions at tree level correspond to the seesaws: the operators $\mathcal{O}_{5,a}^{(0)}$, $\mathcal{O}_{5,c}^{(0)}$ and $\mathcal{O}_{5,b}^{(0)}$ are associated with the Type-I (fermion singlet), Type-II (scalar triplet) and Type-III (fermion triplets) seesaws, respectively. Also, they may be generated at the loop level Ma:2006km (see Ref. Cai:2017jrq for a review of radiative neutrino mass models). Note that there exists only one independent contraction among the fields in the Weinberg operator, though every possible contraction can be obtained from the others via a Fierz transformation Fierz:1937wjm . Indeed, using $\left(\sigma_{i}\right)_{ab}\left(\sigma_{i}\right)_{cd}=2\delta_{ad}\delta_{bc}-\delta_{ab}\delta_{cd}$, with $\sigma_{i}$ ($i=1,2,3$) the Pauli matrices (generators of the SU(2) group) and $(a,b,c,d)$ the $SU(2)$ indices, the different operators in Eq. (6) are related as
@@
In the following, we will focus on the case in which $C_{5,a}^{(0)}$ is very suppressed, so that its contribution to neutrino mass is irrelevant. Also, to alleviate notation, we will denote by $\mathcal{O}^{(0)}_{5}$ the fermion singlet Lorentz contraction, i.e., equivalent to $\mathcal{O}^{(0)}_{5,a}$.
@@ 2.2 Extensions with one new scalar multiplet
Let us discuss first the $\mathcal{O}_{5}^{(1)}$ term. In order to determine the $SU(2)$ representations for $\Phi_{1}$, we first observe the transformation $H\,L_{\alpha}\,L_{\beta}\sim\bf 2\otimes 2\otimes 2=4\oplus 2\oplus 2^{\prime}$ under SU(2). However, if $\Phi_{1}$ is an SU(2) doublet, a two-Higgs-Doublet Model Lee:1973iz (see Ref. Branco:2011iw for a review) is obtained, which has already been widely studied in literature Oliver:2001eg ; Hernandez-Garcia:2019uof . Therefore, $\Phi_{1}\sim{\bf 4}$ is the only interesting possibility for our purposes. The hypercharge of the new scalar quadruplet is fixed by the Weinberg operator and must be either ±$\pm 1/2$ or ±3/2plus-or-minus32$\pm$ 3/2± 3 / 2.111We take the negative hypercharge assignment. Multiplets with positive hypercharges are still a viable option, but a difference in the hypercharge sign does not affect our analysis; therefore we do not consider them without loss of generality.
Let us now consider $\mathcal{O}_{5}^{(2)}$. Whatever the representation of $\Phi_{1}$ is, say $\mathbf{N_{1}}$, it is always possible to construct a singlet with two identical scalar fields and two lepton doublets. However, in this case, the hypercharge of the new multiplet is forced to be ±$\pm 1/2$. Thus, in order to have a neutral scalar component, $\Phi_{1}$ must be in an even SU(2) representation. If we restrict our study to $\mathbf{N_{1}}\leq\mathbf{5}$ in order to avoid problems with unitarity and non-perturbativity close to the EW scale, due to their large RGE running Hally:2012pu ; Earl:2013jsa , the only viable representations are the doublet (which corresponds to the already mentioned 2HDM) and, again, the quadruplet. Neglecting the 2HDM-like models, the only viable possibilities in $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ are the following:
@@ 2.3 Extensions with two new scalar multiplets
Figure 3: Feynman diagram associated with the new Weinberg-like operators $\mathcal{O}^{(3)}_{5}$, which may arise due to the presence of two new scalar multiplets, $\Phi_{1}$ and $\Phi_{2}$.
@@
In this framework, given the two scalar multiplets representations $\mathbf{N_{1}}$ and $\mathbf{N_{2}}$, we need to construct a singlet among $\Phi_{1}$, $\Phi_{2}$ and two lepton doublets. Using $SU(2)$ properties, we require either $\mathbf{N_{1}}\otimes\mathbf{N_{2}}=\mathbf{1}$ or $\mathbf{N_{1}}\otimes\mathbf{N_{2}}=\mathbf{3}$. The former is possible only for $\mathbf{N_{1}}=\mathbf{N_{2}}$, whereas, in the latter case, we have $\mathbf{N_{1}}\otimes\mathbf{3}=\mathbf{N_{1}-2}\oplus\mathbf{N_{1}}\oplus\mathbf{N_{1}+2}$. Therefore, in addition to the case of being equal, $\mathbf{N_{1}}$ and $\mathbf{N_{2}}$ may be either two consecutive even or two consecutive odd representations. In order to construct the Weinberg-like operator $\mathcal{O}_{5}^{(3)}$, we also require that the hypercharges of $\Phi_{1}$ and $\Phi_{2}$ are either equal in modulus or differ by one unit. As before, we consider representations $\mathbf{N_{1,2}}\leq\mathbf{5}$, neglect scenarios containing doublets or two identical representations with identical hypercharges and require that each scalar multiplet contains a neutral component. This leaves us with the following six possibilities:
@@
Once again, a change in sign for the hypercharge of each scalar multiplet will not affect the results of our analyses; thus, one can also write these scenarios with positive hypercharges. Further, it can be seen that the $\mathbf{B_{I}}$ scenario contains multiplets from the class-A scenarios, hence, in this case neutrino masses may also obtain a contribution from $\mathcal{O}_{5}^{(1,2)}$ apart from $\mathcal{O}_{5}^{(3)}$. Finally, it is worth mentioning that in the $\mathbf{B_{V,VI}}$ scenarios containing a scalar triplet with hypercharge $-1$, a Type-II seesaw contribution may be induced. In these cases, in order to have $C^{(3)}_{5}\gtrsim C^{(0)}_{5}$, the trilinear coupling of the triplet with the Higgs doublets would have to be suppressed.
@@ 2.4 Wilson Coefficients for neutrino masses
The representation $\mathbf{N}$ of the fermion mediator can be determined from the the representation of the two scalar multiplets ($\mathbf{N_{1}}$ and $\mathbf{N_{2}}$) entering a particular Weinberg operator (including the SM Higgs doublet). The contraction $\left(\overline{L}\Phi_{i}\right)$ gives $\mathbf{N_{i}}\otimes 2=(\mathbf{N_{i}-1})\oplus(\mathbf{N_{i}+1})$. Thus, if the representations of the two scalar multiplets are identical, the possible contractions of the full Weinberg operator are either from two $\mathbf{(N_{i}-1)}$-tuplets or from two $\mathbf{(N_{i}+1)}$-tuplets, hence $\mathbf{N}=\mathbf{N_{i}\pm 1}$. On the other hand, if $\mathbf{N_{1}}\neq\mathbf{N_{2}}$, for instance, when $\mathbf{N_{1}}$ and $\mathbf{N_{2}}$ are two consecutive odd or even representations, the only possible contraction is between two $(\mathbf{N_{1}+1})$-tuplets, where $\mathbf{N_{1}}$ is the smaller of the two representations, thus $\mathbf{N}=\mathbf{N_{1}}+1$ for $\mathbf{N_{1}}<\mathbf{N_{2}}$.
@@ 3.1 The scalar potential
$V^{\mathbf{A_{II}}}_{L}(H,\Phi_{1})=$ $\;\,\mu_{H}^{2}(H^{\dagger}H)+\mu_{\Phi_{1}}\Phi_{1}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{1}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\lambda_{2}(H^{\dagger}H)^{2}+\lambda_{3}(H^{\dagger}H)(\Phi_{1}^{\dagger}\Phi_{1})+$ (13)
@@
$\mathbf{B_{V}}$ $\mu_{1}\Phi_{1}\Phi_{2}^{\ast}\Phi_{2}+\mu_{2}\Phi_{1}H^{\ast}H+\mu_{3}\Phi_{2}HH+$
$+\lambda_{2}\Phi_{1}\Phi_{1}\Phi_{1}\Phi_{1}+\lambda_{3}H^{\ast}H\Phi_{1}\Phi_{1}+{\rm H.c.}$
$\mathbf{B_{VI}}$ $\mu_{1}\Phi_{1}\Phi_{1}^{\ast}\Phi_{2}+\mu_{2}\Phi_{1}HH+\lambda_{2}\Phi_{2}\Phi_{2}\Phi_{1}^{\ast}\Phi_{1}+$
$+\lambda_{3}H^{\ast}H\Phi_{2}\Phi_{2}+\lambda_{4}\Phi_{2}\Phi_{2}\Phi_{2}\Phi_{2}+{\rm H.c.}$
Table 3: Scalar potential terms violating $U(1)_{L}$ and $U(1)_{X}$ accidental symmetries in the class-B scenarios. We denote the term $\lambda_{3}$ in $\mathbf{B_{I}}$ by a ${}^{\prime}$ as there are two independent contractions associated with the fields involved in the operator.
@@
Let us start with the class-A scenarios, in which we add to the SM only one scalar quadruplet with hypercharge $-1/2$ or $-3/2$. Given the products shown above, the only allowed potential term linear in a quadruplet $\Phi_{1}$ is with three SM Higgs. For both hypercharge assignments $-1/2$ and $-3/2$, an invariant mixed potential term of the form $\Phi_{1}H^{3}$ is allowed (see terms with $\lambda_{8}$ and $\lambda_{6}$ in Eqs. (11) and (14), respectively). In the class B scenarios, with quadruplets ($\mathbf{B_{I}}$) we can again have $\Phi_{i}H^{3}$ terms ($\lambda_{5}$ and $\lambda_{6}$ operators in Table 3); with triplets, on the other hand, it is possible to construct invariant terms of the form $\mu\Phi_{i}H^{2}$, where $\mu$ is some dimensionful coupling. This is the case of $\mathbf{B_{II,V,VI}}$ scenarios. SU(2) quintuplets do not allow for linear terms.
However, there is an interesting scalar potential operator which can always be constructed in class-B scenarios, the mixed linear term $\lambda_{1}\Phi_{1}\Phi_{2}H^{2}$. If the two fields transform under the same representation, the product $\Phi_{1}\Phi_{2}$ contains a singlet and the new mixed scalar term can always be written. If the two multiplets are on the other hand in two consecutive even representations, $\mathbf{N_{1}=2a}$ and $\mathbf{N_{2}=2(a+1)}$ with $\mathbf{a}\geq 1$, the product among the two representations is
$\mathbf{2a}\otimes\mathbf{2(a+1)}=$ $\mathbf{3}\oplus\mathbf{5}\oplus...\oplus\mathbf{4a+}1\,,$ (17)
and it is clear that from their product that we can always build a triplet. Finally, in the case of two odd consecutive representations, i.e., $\mathbf{N_{1}=2a+1}$ and $\mathbf{N_{2}=2a+3}$, we analogously have
$\mathbf{2a+1}\otimes\mathbf{2a+3}=$ $\mathbf{3}\oplus\mathbf{5}\oplus...\oplus\mathbf{4a+3}\,,$ (18)
where, once again, we obtain the same result. Thus, for all the $\mathbf{B}$ class scenarios, the mixed term is always allowed.

### v2

@@ Abstract
Lepton Number Violation, Multi-Higgs Models
@@ 1 Introduction
Considering that $C^{(i)}_{n}\equiv c^{(i)}_{n}/\Lambda$, with $c^{(i)}_{n}$ a dimensionless parameter and $\Lambda$ the new physics scale, Eq. (3) yields an explanation for the smallness of neutrino masses, which scale as $v^{2}/\Lambda$, where $v$ is the Vacuum Expectation Value (VEV) of the SM Higgs field, $\left\langle H\right\rangle\equiv v=174$ GeV. For order one couplings, $c^{(0)}_{5,a}\sim\mathcal{O}(1)$, reproducing the atmospheric mass scale, $m^{(0)}_{\nu}\gtrsim\mathcal{O}(0.05)$ eV, demands $\Lambda\lesssim 10^{14}$ GeV, close to the scale of Grand Unified Theories. This is the seesaw mechanism Yanagida:1980xy . If this is the origin of neutrino masses, it will therefore be very difficult to probe the new physics responsible for lepton number violation.
@@
The possible new Weinberg-like operators with up to two new scalars $\Phi_{1},\Phi_{2}$ are
@@
These have to be understood as classes of operators, encoding all the allowed possible conjugations of the scalar fields. For instance, ${\mathcal{O}}^{(1)}_{5}$ may involve $\widetilde{H}$ instead of $H$ [^1] Note that, if considered individually, operators with for example $\Phi^{c}_{1}$ instead of $\Phi_{1}$ would correspond to just a redefinition of the new field. However, in class-$\bf B$ scenarios (see later) more than one operator is generated and therefore whether the field or its conjugate appears is relevant.
@@
In this work, we study and provide a catalogue of the new scenarios generated by these new Weinberg-like operators constructed with extra scalar multiplets that take a VEV. We consider natural scenarios, in which the new scalar multiplets obtain suppressed induced VEVs, i.e., $v_{i}\sim v^{3}/m_{\Phi}^{2}$. The most appealing scenarios are the ones in which the dominant contribution to neutrino masses is provided by the new operators. The UV models that generate those at tree level are considered genuine and will be studied in a companion paper Giarnetti:2023osf .
@@ 2.1 Neutrino masses in the Standard Model Effective Field Theory
$\mathcal{O}^{(0)}_{5,a}=(HL)_{\mathbf{1}}(HL)_{\mathbf{1}}\,,$ $\mathcal{O}^{(0)}_{5,b}=(HL)_{\mathbf{3}}(HL)_{\mathbf{3}}$ (6)
$\mathcal{O}^{(0)}_{5,c}=(HH)_{\mathbf{3}}(LL)_{\mathbf{3}}\,,$ $\mathcal{O}^{(0)}_{5,d}=(HH)_{\mathbf{1}}(LL)_{\mathbf{1}}.$
Since the singlet $(HH)_{1}=0$, the last operator identically vanishes. The other cases yield non-zero neutrino masses and their UV completions at tree level correspond to the seesaws: the operators $\mathcal{O}_{5,a}^{(0)}$, $\mathcal{O}_{5,c}^{(0)}$ and $\mathcal{O}_{5,b}^{(0)}$ are associated with the Type-I (fermion singlet), Type-II (scalar triplet) and Type-III (fermion triplet) seesaws, respectively. Also, they may be generated at the loop level Ma:2006km (see Ref. Cai:2017jrq for a review of radiative neutrino mass models). Note that there exists only one independent contraction among the fields in the Weinberg operator, though every possible contraction can be obtained from the others via a Fierz transformation Fierz:1937wjm . Indeed, using $\left(\sigma_{i}\right)_{ab}\left(\sigma_{i}\right)_{cd}=2\delta_{ad}\delta_{bc}-\delta_{ab}\delta_{cd}$, with $\sigma_{i}$ ($i=1,2,3$) the Pauli matrices (generators of the $SU(2)$ group) and $(a,b,c,d)$ the $SU(2)$ indices, the different operators in Eq. (6) are related as
@@
In the following, we will focus on the case in which $C_{5,a}^{(0)}$ is very suppressed, so that its contribution to neutrino mass is irrelevant. Also, to alleviate notation, we will denote the fermion singlet Lorentz contraction ($\mathcal{O}^{(0)}_{5,a}$) by $\mathcal{O}^{(0)}_{5}$.
@@ 2.2 Extensions with one new scalar multiplet
Let us discuss first the $\mathcal{O}_{5}^{(1)}$ term. In order to determine the $SU(2)$ representations for $\Phi_{1}$, we first observe the transformation $H\,L_{\alpha}\,L_{\beta}\sim\bf 2\otimes 2\otimes 2=4\oplus 2\oplus 2^{\prime}$ under SU(2). However, if $\Phi_{1}$ is an $SU(2)$ doublet, a two-Higgs-Doublet Model Lee:1973iz (see Ref. Branco:2011iw for a review) is obtained, which has already been widely studied in literature Oliver:2001eg ; Hernandez-Garcia:2019uof . Therefore, $\Phi_{1}\sim{\bf 4}$ is the only interesting possibility for our purposes. The hypercharge of the new scalar quadruplet is fixed by the Weinberg operator, and we take the negative hypercharge assignment without loss of generality, $-1/2$ or 3/232-3/2- 3 / 2.222Note that both positive and negative hypercharge assignments, ±$\pm 1/2$ or ±$\pm 3/2$, are allowed, depending on the definition of the effective operator: $\Phi_{1}\sim(4,-1/2)$ for $\overline{L}\Phi_{1}\widetilde{H}^{T}L^{c}$ or $\Phi_{1}\sim(4,+1/2)$ for $\overline{L}\Phi_{1}^{c}\widetilde{H}^{T}L^{c}$; similarly, $\Phi_{1}\sim(4,-3/2)$ for $\overline{L}\Phi_{1}H^{T}L^{c}$, or $\Phi_{1}\sim(4,+3/2)$ for $\overline{L}{\Phi}^{c}_{1}H^{T}L^{c}$. Let us now consider $\mathcal{O}_{5}^{(2)}$. Whatever the representation of $\Phi_{1}$ is, say $\mathbf{N_{1}}$, it is always possible to construct a singlet with two identical scalar fields and two lepton doublets. However, in this case, the hypercharge of the new multiplet is forced to be $-1/2$. Thus, in order to have a neutral scalar component, $\Phi_{1}$ must be in an even $SU(2)$ representation. If we restrict our study to $\mathbf{N_{1}}\leq\mathbf{5}$ in order to avoid problems with unitarity and non-perturbativity close to the EW scale, due to their large RGE running Hally:2012pu ; Earl:2013jsa , the only viable representations are the doublet (which corresponds to the already mentioned 2HDM) and, again, the quadruplet. Neglecting the 2HDM-like models, the only viable possibilities in $\mathcal{O}_{5}^{(1)}$ and $\mathcal{O}_{5}^{(2)}$ are the following:
@@ 2.3 Extensions with two new scalar multiplets
Figure 3: Feynman diagram associated with the new Weinberg-like operator $\mathcal{O}^{(3)}_{5}$, which may arise due to the presence of two new scalar multiplets, $\Phi_{1}$ and $\Phi_{2}$.
@@
[…]
@@
Once again, a change in sign for the hypercharge of each scalar multiplet will not affect the results of our analyses; thus, one can also write these scenarios with opposite hypercharges. Notice that, given the cut on the $\Phi_{i}$ representations, there are no scenarios where the two scalar multiplets have the same hypercharge. Further, it can be seen that the $\mathbf{B_{I}}$ scenario contains multiplets from the class-A scenarios, hence, in this case neutrino masses may also obtain a contribution from $\mathcal{O}_{5}^{(1,2)}$ apart from $\mathcal{O}_{5}^{(3)}$. Finally, it is worth mentioning that in the $\mathbf{B_{V,VI}}$ scenarios containing a scalar triplet with hypercharge $-1$, a Type-II seesaw contribution may be induced. In these cases, in order to have $C^{(3)}_{5}\gtrsim C^{(0)}_{5}$, the trilinear coupling of the triplet with the Higgs doublets would have to be suppressed. Despite being in principle non-genuine, we will consider also these two scenarios in the rest of the paper under the hypothesis that $C^{(0)}_{5}$ is negligible compared to $C^{(3)}_{5}$.
@@ 2.4 Wilson Coefficients for neutrino masses
The representation $\mathbf{N}$ of the fermion mediator can be determined from the representation of the two scalar multiplets ($\mathbf{N_{1}}$ and $\mathbf{N_{2}}$) entering a particular Weinberg operator (including the SM Higgs doublet). The contraction $\left(\overline{L}\Phi_{i}\right)$ gives $\mathbf{N_{i}}\otimes\mathbf{2}=(\mathbf{N_{i}-1})\oplus(\mathbf{N_{i}+1})$. Thus, if the representations of the two scalar multiplets are identical, the possible contractions of the full Weinberg operator are either from two $\mathbf{(N_{i}-1)}$-tuplets or from two $\mathbf{(N_{i}+1)}$-tuplets, hence $\mathbf{N}=\mathbf{N_{i}\pm 1}$. On the other hand, if $\mathbf{N_{1}}\neq\mathbf{N_{2}}$, for instance, when $\mathbf{N_{1}}$ and $\mathbf{N_{2}}$ are two consecutive odd or even representations, the only possible contraction is between two $(\mathbf{N_{1}+1})$-tuplets, where $\mathbf{N_{1}}$ is the smaller of the two representations, thus $\mathbf{N}=\mathbf{N_{1}}+1$ for $\mathbf{N_{1}}<\mathbf{N_{2}}$.
@@ 3.1 The scalar potential
$V^{\mathbf{A_{II}}}_{L}(H,\Phi_{1})=$ $\;\,\mu_{H}^{2}(H^{\dagger}H)+\mu_{\Phi_{1}}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{1}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\lambda_{2}(H^{\dagger}H)^{2}+\lambda_{3}(H^{\dagger}H)(\Phi_{1}^{\dagger}\Phi_{1})+$ (13)
@@
$\mathbf{B_{V}}$ $\mu_{1}\Phi_{1}\Phi_{2}^{\ast}\Phi_{2}+\mu_{2}\Phi_{1}H^{\ast}H+\mu_{3}\Phi_{2}HH+{\mu_{4}^{2}\,\Phi_{1}^{2}}$
$+\lambda_{2}\left(\Phi_{1}\Phi_{1}\Phi_{1}\Phi_{1}\right)+\lambda_{3}\left(H^{\ast}H\Phi_{1}\Phi_{1}\right)+{\lambda_{4}\,\Phi_{1}^{2}\Phi_{2}^{\ast}\Phi_{2}}+{\rm H.c.}$
$\mathbf{B_{VI}}$ $\mu_{1}\Phi_{1}\Phi_{1}^{\ast}\Phi_{2}+\mu_{2}\Phi_{1}HH+{\mu_{3}^{2}\,\Phi_{2}^{2}}+{\mu_{4}\,\Phi_{2}^{3}}+$
$+\lambda_{2}\left(\Phi_{2}\Phi_{2}\Phi_{1}^{\ast}\Phi_{1}\right)+\lambda_{3}\left(H^{\ast}H\Phi_{2}\Phi_{2}\right)+\lambda_{4}\Phi_{2}\Phi_{2}\Phi_{2}\Phi_{2}+{\rm H.c.}$
Table 3: Scalar potential terms violating $U(1)_{L}$ and $U(1)_{X}$ accidental symmetries in the class-B scenarios. We denote the term $\lambda_{3}$ in $\mathbf{B_{I}}$ by a as there are two independent contractions associated with the fields involved in the operator.
@@
Let us start with the class-A scenarios, in which we add to the SM only one scalar quadruplet with hypercharge $-1/2$ or $-3/2$. Given the products shown above, the only allowed potential term linear in a quadruplet $\Phi_{1}$ is with three SM Higgs. For both hypercharge assignments $-1/2$ and $-3/2$, an invariant mixed potential term of the form $\Phi_{1}H^{3}$ is allowed (see terms with $\lambda_{8}$ and $\lambda_{6}$ in Eqs. (11) and (14), respectively). In the class B scenarios, with quadruplets ($\mathbf{B_{I}}$) we can again have $\Phi_{i}H^{3}$ terms ($\lambda_{5}$ and $\lambda_{6}$ terms in Table 3); with triplets, on the other hand, it is possible to construct invariant terms of the form $\mu\Phi_{i}H^{2}$, where $\mu$ is some dimensionful coupling. This is the case of $\mathbf{B_{II,V,VI}}$ scenarios. $SU(2)$ quintuplets do not allow for linear terms.
However, there is an interesting scalar potential operator which can always be constructed in class-B scenarios, the mixed linear term $\lambda_{1}\Phi_{1}\Phi_{2}H^{2}$. Indeed, as shown in Sec. 2.4, the contraction $(LL)_{3}(\Phi_{1}\Phi_{2})_{3}$ is always allowed in our scenarios. Since the combination $(HH)$ has the same $SU(2)$ structure of the combination $(LL)$, the scalar potential must contain a term coming from the contraction $(HH)_{3}(\Phi_{1}\Phi_{2})_{3}$.